OpenStudy (anonymous):

25. If ten resistances each of 10 ohms are connected in parallel then the total resistance is A) 1 ohm B) 5 ohms C) 2 ohms D) 10 ohms

OpenStudy (souvik):

let the total resistance be R do 1/R=1/10+1/10+1/10+...10 terms or 1/R=10/10 or R= 1 ohm

OpenStudy (anonymous):

thats what i got but their answer is c